OpenStudy (anonymous):

solve for x. x-1/4=51/5 x=

OpenStudy (akashdeepdeb):

\[x - \frac{1}{4} = \frac{51}{5}\] Add \(\frac{1}{4}\) to both sides. \[x = \frac{51}{5} + \frac{1}{4}\] Find the LCM, 5 and 4's LCM is = 20. \[x = \frac{51*4 + 1*5}{20}\] \[x = \frac{209}{20}\] Getting this? :)

jhonyy9 (jhonyy9):

i think there wann being (x-1) 51 ----- = --- so then cross multiplie and will get 4 5 5(x-1)= 4*51 5x-5 = 204 5x = 209 x = 209/5 hope this is understandably

OpenStudy (anonymous):

thank you!